OpenStudy (anonymous):

S=2lw+2wh+2hl S=800 l=10 h=20 800=2(10w)+2w(20)+(2)(20)(10) 800=60w+400 400=60w w = 6.6 is it right?

OpenStudy (ahsome):

@5454354366464 Are you still here?

OpenStudy (anonymous):

yes

OpenStudy (ahsome):

Good Equation=\[S=2(lw+wh+hl)\]\[S=800\]\[l=10\]\[w=w\]\[h=20\] Sub values: \[S=2(lw+wh+hl)\]\[800=2(10w+20w+20*10)\]\[800=2(10w+20w+200)\]\[800=20w+40w+400\]\[400=20w+40w\]\[400=60w\]\[w=\frac{400}{60}\]\[w=\frac{20}{3}\]\[w=6.666666...\] So Yes, you were right :)

OpenStudy (anonymous):

thanks'